OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10) A. y = x B. y = x + 6 C. y = x – 6 D. y = x + 6

OpenStudy (anonymous):

Parallel lines only differ by a constant term. So ,let the equation be y = x+ k where k is the new constant. Plug in the point (4, 10) , 10 = 4 + k , k = -6. So the equation is C. y = x – 6 .